An air-dry soil with a moisture content of 35% by weight is to be used in a pot

experiment. If each pot is to have 20 kg oven-dried soil, how much of the air-dry soil should

be weighed into the pot and how much is the moisture of that soil in kg?


1
Expert's answer
2021-11-24T12:22:14-0500

"m_w" =water mass

"m_t" = total mass

"(20+m_w) = m_t \\

m_w = 0.35m_t \\

20 + 0.35m_t = m_t \\

20 = 0.65m_t \\

m_t = 30 .77 \;kg"

30.77 kg of the air-dry soil should be weighed into the pot.

"m_w = 0.35 \times 30.77= 10. 77 \;kg"

10.77 kg of that is moist